OpenStudy (anonymous):

what is the given amount converted to the given unit ? 195 s; minutes

hero (hero):

\[195 \space\text{seconds} \times \frac{\space\text{minutes}}{60 \space\text{seconds}}\]

OpenStudy (anonymous):

So I would multiply 195 and 60 ?

OpenStudy (anonymous):

no I would divide and get 3.25

hero (hero):

3.25 minutes Correct

OpenStudy (anonymous):

Thank u ! could u check another for me ?

OpenStudy (anonymous):

??

hero (hero):

What is it?

hero (hero):

@mikaa_toxica13

OpenStudy (anonymous):

108 ft; inches 9 inches

hero (hero):

You think that 108 ft = 9 inches?

hero (hero):

This is why they teach dimensional analysis\conversion. \[108\space\text{ft} \times \frac{12\space\text{in}}{\text{ft}} = (108)(12)\space\text{in}\]

OpenStudy (anonymous):

Uhmmm I don't think that came out how u wanted it too. I did the same thing I did on other problem and got 9 inches. It's one of the answers provided as well.

hero (hero):

But you should know 108 doesn't equal 9 inches. You can get 9 inches on a little pocket ruler. But to get 108 ft, it would take you a while to measure that using a ruler.